OpenStudy (anonymous):

how do i solve for the variable 3x (2x+1) (x-1) = 0

OpenStudy (anonymous):

this is only 3 possible answers: 3x=0, so x=0 2x+1=0, so x=-1/2 x-1=0, so x=1

OpenStudy (anonymous):

thx